A Ch3Ch2Ch(OH)Ph can be prepared from
1: PhCOCH3+CH3MgBr
2: PhCHO+C2H5MgBr
3: C2H5CHO+PhMgBr
4: All of these
1
Expert's answer
2020-07-27T05:35:01-0400

Here two options are correct. 2nd and 3rd option

From 2nd option -

PhCH=O + C2H5MgBr = PhCH(OH)CH2CH3

From 3rd option

C2H5CHO + PhMgBr = C2H5CH(OH)Ph

In both options grignard reagent will react and after hydrolysis the resultant product will formed.